OpenStudy (anonymous):

Will give Medals!!!! Gordon is evaluating some stocks he wants to purchase. You can see the function of the stock price below. The variable x represents days.

OpenStudy (anonymous):

Stock A B C Price f(x) = 25(1.08)x x f(x) 0 22.00 1 21.56 2 21.13 The initial price of this stock is $30, but it has been increasing 4% each day.

OpenStudy (anonymous):

1. Create the price function for Stock D. It should have the second lowest starting price and the second highest rate of growth. Using complete sentences, justify that your function meets these requirements.

OpenStudy (anonymous):

I'm sorry but i don't really understand the question.

OpenStudy (anonymous):

Price = (starting value) (growth factor) (day number) D = (number between 21.13 and 21.56)((1 +second highest growth factor ) I don't understand either what you are to do with the stock mentioned at end. which would follow price = 30(1.04) x

OpenStudy (anonymous):

maybe its , price = 30(1.04) * f(x) + g(x)